OpenStudy (king):

If motion of a particle given below is x=u(t-2s)+a(t-2s)^2 A)initial velocity is u B)acceleration is a C)acceleration is 2a D)at t=2s particle is at origin More than 1 option can be right Please explain ure answer.

OpenStudy (king):

s is seconds just a unit

OpenStudy (anonymous):

A,C and D are true :-) Compare it with \(x = u\times t + \frac{1}{2}\times a \times t^2\)

OpenStudy (king):

actually ure wrong..i wont tell u the rite answer try it!!

OpenStudy (anonymous):

I can't be wrong

OpenStudy (king):

and please explain...

OpenStudy (king):

no u hav to expand then compare

OpenStudy (anonymous):

Na I don't have to expand since (t - 2) is time only (2 is in seconds and so is t)

OpenStudy (king):

but its not a(t-2s)^2 its supposed a/2

OpenStudy (king):

it at^2 ok but shouldnt it be (at^2)/2

OpenStudy (anonymous):

\[\frac{1}{2}\times 2a\times (t - 2)^2\]

OpenStudy (king):

initial velocity is not u C),D) is rite

OpenStudy (anonymous):

Initial velocity must be U!

OpenStudy (king):

its not..my sir explained but i frgt

OpenStudy (king):

pls expand once aand see what u get

OpenStudy (anonymous):

Okay I get what you meant

OpenStudy (king):

ahh now xplain

OpenStudy (anonymous):

Hmm because it's (t-2)^2 not t^2 so an extra co-efficient of t is there

OpenStudy (king):

oh yeah and btw my sir took x as xfinal - xinitial

OpenStudy (anonymous):

Okay!

OpenStudy (king):

hey explain!